# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v3

@@ Abstract.
Three components: input task, design space, and resulting workflow. Input task is shortening text to specified lengths while preserving meaning. Design space flows through 3 main steps 1) Objectives which are to have a creative response while maintaining accurate actuality, 2) strategies, which are for response diversity, user guidance, and validation, and 3) tactics of: parallel generation, step to select focus, transform outputs, and evaluate candidates. After design space comes the resulting workflow. This workflow inputs a long text, finds areas to edit, fixes those areas, verifies those fixes, and outputs a shorter text with a length slider. The fix and verify steps have parallel generation.
@@ 1. Introduction
People use Large Language Models (LLMs) for assistance with both single-step tasks, like suggesting synonyms, and more complex multi-step tasks, like writing stories, editing text, and organizing ideas. Despite their widespread adoption, LLMs suffer from quality deficits like hallucinations (Dziri et al., 2022; Nair et al., 2023b; Maynez et al., 2020), brittleness to prompt changes (Gero et al., 2023; Arora et al., 2022; Zamfirescu-Pereira et al., 2023; Nie et al., 2022; Holtzman et al., 2021), and user interventions limited to prompting (Wu et al., 2023c; Parameswaran et al., 2023). These deficits are exacerbated for complex tasks, which require multiple steps in which errors can arise and propagate (Dziri et al., 2023). Unfortunately, many common use cases require multiple reasoning steps. For example, the use case of shortening text requires first deciding which portions of the text to edit and then determining how to revise each one (Bernstein et al., 2010).
@@
To tackle complex tasks, recent research has turned to LLM chaining techniques. Chaining decomposes a task into multiple calls to an LLM, in which the output of one call affects the input to the next call (Wu et al., 2021). For example, when shortening text, an LLM chain could identify verbose sentences, edit each one, and propose outputs of variable lengths by composing multiple edits. Chaining has many benefits: this strategy can enable new abilities like long-form story writing (Yang et al., 2022; Mirowski et al., 2023), increase accuracy on logical reasoning questions (Xie et al., 2023; Zhang et al., 2022; Creswell and Shanahan, 2022), and enable greater transparency and debuggability (Huang et al., 2023b; Wu et al., 2022a; Reppert et al., 2023). Although initial evidence suggests that chaining can help mitigate LLM quality deficits, current solutions are imperfect (Nair et al., 2023a; Mirowski et al., 2023; Creswell and Shanahan, 2022; Park et al., 2022; Long et al., 2023; Josifoski et al., 2023). LLM chaining is nascent and challenging, with a seemingly intractable design space.
@@
The challenge of decomposing complex work, however, is not new. In the more established field of crowdsourcing, researchers have faced a similar problem setup. Groups of people (requesters) outsource tasks to other people (crowdworkers) on online platforms. Crowdworkers have their own limitations: approximately 30% of crowdworkers’ outputs are low quality (Bernstein et al., 2010), small changes in instructions affect crowdworker responses (Kittur et al., 2012; McDonnell et al., 2016), and requesters have limited modes of interaction with crowdworkers (Salehi et al., 2017). To compensate for these limitations, crowdsourcing researchers have developed a variety of structured workflows. Crowdsourcing workflows decompose a task into a series of smaller microtasks for crowdworkers to complete. These workflows account for crowdworker limitations by reducing task complexity and leveraging redundancy, aggregation, and explicit validation (Bernstein et al., 2010; Cheng and Bernstein, 2015; Ambati et al., 2012; Zaidan and Callison-Burch, 2011).
@@
Given the parallels between these fields, we investigate how crowdsourcing techniques may be adapted to design LLM chains (Figure 1). Selecting relevant strategies from over a decade of crowdsourcing research is challenging, especially as LLMs and crowdworkers have salient differences (Wu et al., 2023c). For example, crowdworkers may be slow to generate large text amounts, whereas LLMs may lack common sense and factual grounding, making them more sensitive to instructions and likely to “hallucinate” information.
@@
From this review, we also surface the strategies used to marshal and combine tactics to achieve objectives. For example, the objective to encourage creativity in outputs can be achieved by strategies that produce diverse responses, allow user guidance, and utilize adaptable architectures. We gather from the literature how tactics have been used in prior work to support strategies. For example, workflows and chains encourage diverse responses with redundant or communicative tactical architectures. Strategies constrain the space and evaluation of tactic choices for both crowds and LLMs (Figure 4).
To better assess the impact of similarities and differences between crowdworkers and LLMs on the implementation of strategies, we implement three case studies adapting crowdsourcing workflows to use LLMs. We chose workflows that enable direct manipulation of the output, an instance of the user guidance strategy for quality control that has been explored in crowdsourcing, but is underexplored in chaining. The case studies implement tasks that vary in their degrees of creativity, adapting Cascade (Chilton et al., 2013) to control the precision of a generated taxonomy, Soylent (Bernstein et al., 2010) to shorten text, and Mechanical Novel (Kim et al., 2017) to control the narrative elements of a fictional story.
@@
Informed by both our design space and the three case studies, we discuss how differences between LLMs and crowdworkers impact effective tactic designs for LLM chains. For example, because LLM stochasticity differs from human diversity, tactic choices for LLMs must induce response diversity more explicitly and account for limitations in LLM self-validation. We find that our strategies help focus workflow adaptation and suggest areas of promising future work. Such areas include how the relative speed and reduced cost of LLMs relative to crowdworkers presents an opportunity to create higher quality chains, and we recommend improved tools and methods to augment this design process.
2. Related work
The bulk of our design space contribution is a literature review, so most related work has been discussed inline throughout the paper. Concurrent to this paper, a few other groups investigate the intersection of crowdsourcing and LLMs. One paper explores the impact of incorporating LLMs into crowdsourcing workflows on requester and crowdworker stakeholders (ALLEN et al., 2023). A second considers hybrid Crowd-LLM data annotation (Li, 2024). Another implements three strategies from crowdsourcing using LLMs for data processing tasks (Parameswaran et al., 2023). Closest to our work, the final paper has students implement six crowdsourcing workflows. The paper discusses the effectiveness of these workflows and the impact of the difference between LLM and crowdworker abilities on recommendations for chaining (Wu et al., 2023c). We build on this approach of providing recommendations based on case study implementations. We widen the scope of our recommendations by first integrating a broad survey of LLM chaining and crowdsourcing workflow literature into a design space and then using the case studies to explore open questions in the design space. This broader investigation enables us to build a design space and informs more comprehensive recommendations for navigating design decisions and for orienting future work. These recommendations span from higher-level guidance, such as adhering to the tradeoff between outcome quality and resource constraints, to lower-level suggestions, such as pursuing architecture-based methods for quality validation.
@@ 3. Design space methodology
We survey papers that involve the design and implementation of crowdsourcing workflows or LLM chains. Although crowdsourcing workflows and LLM chains have many overlapping concepts, the terms used to denote these concepts differ. In this paper, we refer to both LLMs and crowdworkers as “workers.” We use the term “workflow” to refer to both crowdsourcing workflows and LLM chains. We refer to intermediate steps of workflows or chains as “subtasks.” We use the term “user” to describe the person receiving workflow outputs (e.g., the “requester” in crowdsourcing terms).
@@ 3.1. Paper collection
Workflow literature. We collect papers through search of the keywords: crowd(sourcing) workflow, crowd(sourcing) pipeline, crowdsourcing complex work and microtask. We search the ACM CHI Conference on Human Factors in Computing Systems, ACM Conference on Computer-supported Cooperative Work and Social Computing, the ACM Symposium on User Interface Software and Technology, and the AAAI Conference on Human Computation and Crowdsourcing, as well as Semantic Scholar (Fricke, 2018) and Google Scholar.
Chaining literature. We collect papers through search of the keywords: LLM chain(ing), prompt chain(ing), LLM pipeline, and AI chain(s/ing). We search the International Conference on Learning Representations, the Conference on Neural Information Processing Systems, the International Conference on Machine Learning, the Association for Computational Linguistics proceedings, the Association for Computing Machinery proceedings, as well as Semantic Scholar (Fricke, 2018) and Google Scholar. We include preprints in our literature search.
Paper inclusion criteria. We include papers that implement, or support the implementation of, a workflow with at least two steps; i.e., workflows in which the outputs from one worker inform the input to another worker. We include papers that focus on building an interface or domain specific language for other designers to build workflows. We restrict our search to academic papers, rather than software like LangChain (LangChain, 2022). We exclude works that call multiple workers in parallel but never in a chain of steps (e.g., majority voting of labels). We also exclude techniques such as Chain of Thought (Wei et al., 2022) which issue multiple tasks to a worker in one step instead of incorporating multiple worker calls (Cheng et al., 2015). We include instances in which workers synchronously work on the same interface, viewing and editing each other’s outputs. Finally, the task at hand must have a user working towards some end goal, so we do not consider open-ended collaborative projects like Wikipedia. We make one exception for a paper that explicitly draws from crowdsourcing strategies to complete data aggregation tasks with LLMs (Parameswaran et al., 2023). Although they do not implement chains, the connection to crowdsourcing resonates with our design space goals.
@@ 3.2. Coding strategy
We conduct a thematic analysis (Braun and Clarke, 2006), following the methodologies of similar design spaces in HCI (Lai et al., 2023; Bae et al., 2022; Shi et al., 2021). First, one author extracted 1) the outcomes (objectives) and 2) the elements (tactics) of the workflows from the paper set. We then employed iterative open coding on the extracted data, in which two authors separately coded each data extraction, sorted resulting codes into sub-categories, then discussed these categorizations. In three further rounds of iteration, the same authors refined these categorizations. All authors reviewed the resulting design space.
@@ 4. Design Space
Our design space has two parts: the objectives span the space of a designer’s goals, and tactics span the space of workflow implementations. These objectives are 1) promoting outcome qualities while respecting resource constraints, 2) increasing the accuracy and/or creativity of outputs, and defining the generality or specificity of the set of tasks that the designer wants the workflow to support. For the workflow tactics, workers complete various subtasks that are linked together into the workflow’s architecture.
Title of ”Objectives, what the designer wants the workflow to accomplish”. Three rows of text each with two objectives, examples, and a description. Row 1. A tradeoff between outcome quality (ex. transparency, no hallucination) and resource constraints (ex. effort, latency, expense). Description: ”Designers must work within resource constraints to improve outcome quality”. Row 2. A spectrum between accuracy (ex. math, question-answering) and creativity (ex. writing, prototyping). Description: Different tasks may require accuracy, creativity, or some combination thereof. Row 3. A spectrum between task specific (ex. only applied to one problem (e.g., translation workflow)) and task general (ex. can be applied to many problems(e.g., map-reduce)). Description: Workflows vary from supporting a specific task to supporting a general set of tasks.
@@ 4.1. Objectives in the design space
The objectives of a workflow design define the designer’s goals within several dimensions. Designers will design for some point along a outcome quality vs resource constraints tradeoff, in a range from accuracy to creativity, and in a range spanning between task specificity and generality. (Figure 2).
@@ 4.1.1. Outcome quality vs resource constraints
When working towards outcome qualities, designers must be mindful of resource constraints. Workflows can support quality outputs such as by improving accuracy and/or creativity (Kim et al., 2023b; Creswell and Shanahan, 2022; Lasecki et al., 2013; Reppert et al., 2023; Wu et al., 2021), transparency (Creswell et al., 2022; Creswell and Shanahan, 2022; Wu et al., 2023c; Reppert et al., 2023; Wu et al., 2023a; Kim et al., 2023b), debuggability (Wu et al., 2022a, 2021), and modularity (Huang et al., 2023b; LaToza et al., 2014; Lee et al., 2017; Retelny et al., 2014; Zharikova et al., 2023; Yao et al., 2023; Wu et al., 2023a).
@@
Workflows can also address, although imperfectly, quality concerns. Some quality concerns are induced by the workers, such as hallucinations (Nair et al., 2023a; Mirowski et al., 2023; Creswell and Shanahan, 2022; Nair et al., 2023b; Wu et al., 2023a; Parameswaran et al., 2023) and generative loops (Mirowski et al., 2023) in the case of LLMs, and fatigue (Dai et al., 2013; Alshaibani et al., 2020) in the case of human workers. Others are induced by the workflow itself, such as cascading errors (Wu et al., 2022a, 2021; Xie et al., 2023) and lack of coherency (Bernstein et al., 2010; Yang et al., 2022; Kim et al., 2017; Hahn et al., 2016; Mirowski et al., 2023; Wu et al., 2022a; Hahn et al., 2016). More concerns are induced by the broader context, such as the need to scale (Arora et al., 2022; Kulkarni et al., 2014, 2012), the workflow’s placement within a larger system (Pavel et al., 2014; Christoforaki and Ipeirotis, 2014; Cheng and Bernstein, 2015), unethical use of the workflow (Park et al., 2022), and the impact on human labor (Mirowski et al., 2023; Wu et al., 2023a), sense of ownership (Mirowski et al., 2023) and sense of agency (Kim et al., 2023b; Mirowski et al., 2023). However, mechanisms for improving output quality can be more expensive (Yao et al., 2023; Huffaker et al., 2020), take longer (Chang et al., 2017; Gouravajhala et al., 2018), and require more effort to create (Kittur et al., 2011; LaToza et al., 2014).
@@
Expense. The expense of a workflow refers to paying crowdworkers or paying for an API call to an LLM. Workflows are often more expensive than simpler approaches (e.g., Chain of Thought prompting) (Zharikova et al., 2023; Yao et al., 2023; Wang et al., 2022; Du et al., 2023; Gero et al., 2023; Hahn et al., 2016; Huffaker et al., 2020; Lin et al., 2014a; Dai et al., 2010; Marcus et al., 2011a, b; Nair et al., 2023a). However, the cost is often less than hiring professionals (Ambati et al., 2012; Christoforaki and Ipeirotis, 2014; Zaidan and Callison-Burch, 2011; Liem et al., 2011; Fleming et al., 2023), pretraining models (Kim et al., 2023b; Levine et al., 2022; Yao et al., 2023; Huang et al., 2023b; Cobbe et al., 2021), or completing the task yourself (Zhang et al., 2022; Wang et al., 2023).
@@
Latency. Another constraint is the time for the workflow to complete (Lasecki et al., 2015; Kulkarni et al., 2014; Marcus et al., 2011a, b). Many workflows have lower latency than a single person performing the same task, as work can be parallelized across multiple workers (Teevan et al., 2016; Lasecki et al., 2015; Chilton et al., 2014; Guo et al., 2016; Lee et al., 2017; André et al., 2013; Pietrowicz et al., 2013; Lin et al., 2014a; Gouravajhala et al., 2018; Kulkarni et al., 2014).
@@
Effort. Workflows can help users complete tasks with less effort (Mirowski et al., 2023; Wu et al., 2023a; Kulkarni et al., 2014; Long et al., 2023). However, workflows can require high effort to create (Kulkarni et al., 2012; Hahn et al., 2016; Kittur et al., 2012) and oversee (Wu et al., 2022a; LaToza et al., 2014).
4.1.2. Accuracy vs creativity

### v4

@@ Abstract.
Three components: input task, design space, and resulting workflow. Input task is shortening text to specified lengths while preserving meaning. Design space flows through 3 main steps 1) Objectives which are to have a subjective style while maintaining objective factuality, 2) strategies, which are for response diversity, user guidance, and validation, and 3) tactics of: parallel generation, step to select focus, evaluate outputs, and evaluate candidates. After design space comes the resulting workflow. This workflow inputs a long text, finds areas to edit, fixes those areas, verifies those fixes, and outputs a shorter text with a length slider. The fix and verify steps have parallel generation.
@@ 1. Introduction
People use Large Language Models (LLMs) for assistance with both single-step tasks, like suggesting synonyms, and more complex multi-step tasks, like writing stories, editing text, and organizing ideas. Despite their widespread adoption, LLMs suffer from quality deficits like hallucinations (Dziri et al., 2022; Nair et al., 2023b; Maynez et al., 2020), brittleness to prompt changes (Zamfirescu-Pereira et al., 2023; Nie et al., 2022; Holtzman et al., 2021), and user interventions limited to prompting (Wu et al., 2023b; Parameswaran et al., 2023). These deficits are exacerbated for complex tasks, which require multiple steps in which errors can arise and propagate (Dziri et al., 2023). Unfortunately, many common use cases require multiple reasoning steps. For example, the use case of shortening text requires first deciding which portions of the text to edit and then determining how to revise each one (Bernstein et al., 2010).
@@
To tackle complex tasks, recent research has turned to LLM chaining techniques. Chaining decomposes a task into multiple calls to an LLM, in which the output of one call affects the input to the next call (Wu et al., 2021). For example, when shortening text, an LLM chain could identify verbose sentences, edit each one, and propose outputs of variable lengths by composing multiple edits. Chaining has many benefits: this strategy can enable new abilities like long-form story writing (Yang et al., 2022; Mirowski et al., 2023), increase accuracy on logical reasoning questions (Xie et al., 2023; Zhang et al., 2022; Creswell and Shanahan, 2022), and enable greater transparency and debuggability (Huang et al., 2023b; Wu et al., 2022a; Reppert et al., 2023). LLM chaining is a nascent subfield, with seminal papers emerging in only the past few years (Wu et al., 2021; Josifoski et al., 2023; Wu et al., 2023a; Li et al., 2024). Initial evidence suggests that chaining can mitigate quality deficits, but these solutions continue to face problems such as hallucination (Schick et al., 2022), bias (Mirowski et al., 2023), incorrect responses (Nair et al., 2023a), and expensiveness (Zharikova et al., 2023). There is a seemingly intractable design space of possible chaining combinations from the suggested operators of this early work, so realizing the potential of chaining is challenging due to the design space’s size and complexity.
@@
The challenge of decomposing complex work, however, is not new. In the more established field of crowdsourcing, researchers have faced a similar problem setup. Groups of people (requesters) outsource tasks to other people (crowdworkers) on online platforms. Crowdworkers have their own limitations: crowdworkers may output low quality responses (Bernstein et al., 2010; Alshaibani et al., 2020; Butler et al., 2017; Lasecki et al., 2013), small changes in instructions affect crowdworker responses (Kittur et al., 2012; McDonnell et al., 2016), and requesters have limited modes of interaction with crowdworkers (Salehi et al., 2017). To compensate for these limitations, crowdsourcing researchers have developed a variety of structured workflows. Crowdsourcing workflows decompose a task into a series of smaller microtasks for crowdworkers to complete. These workflows account for crowdworker limitations by reducing task complexity and leveraging redundancy, aggregation, and explicit validation (Bernstein et al., 2010; Cheng and Bernstein, 2015; Ambati et al., 2012; Zaidan and Callison-Burch, 2011).
@@
Given the parallels between these fields, we investigate how crowdsourcing techniques may be adapted to design LLM chains (Figure 1). Selecting relevant strategies from over a decade of crowdsourcing research is challenging, especially as LLMs and crowdworkers have salient differences (Wu et al., 2023b). For example, crowdworkers may be slow to generate large text amounts, whereas LLMs may lack common sense and factual grounding, making them more sensitive to instructions and likely to “hallucinate” information.
@@
From this review, we also surface the strategies used to marshal and combine tactics to achieve objectives. For example, the objective to encourage a subjective output quality can be achieved by strategies that produce diverse responses, allow user guidance, and utilize adaptable architectures. We gather from the literature how tactics have been used in prior work to support strategies. For example, workflows and chains encourage diverse responses with redundant or communicative tactical architectures. Strategies constrain the space and evaluation of tactic choices for both crowds and LLMs (Figure 5).
To better assess the impact of similarities and differences between crowdworkers and LLMs on the implementation of strategies, we implement three case studies adapting crowdsourcing workflows to use LLMs. We chose workflows that enable direct manipulation of the output, an instance of the user guidance strategy for quality control that has been explored in crowdsourcing, but is underexplored in chaining. The case studies implement tasks that vary in their degrees of subjectivity, adapting Cascade (Chilton et al., 2013) to control the precision of a generated taxonomy, Soylent (Bernstein et al., 2010) to control the length of shortened text, and Mechanical Novel (Kim et al., 2017) to control the narrative elements of a fictional story.
@@
Informed by both our design space and the three case studies, we discuss how differences between LLMs and crowdworkers impact effective tactic designs for LLM chains. For example, because LLM stochasticity differs from human diversity, tactic choices for LLMs must induce response diversity more explicitly and account for limitations in LLM self-validation. We find that our strategies help focus workflow adaptation and suggest areas of promising future work. Such areas include how the speed and reduced cost of LLMs relative to crowdworkers presents an opportunity to create higher quality chains, and we recommend improved tools and methods to augment this design process.
2. Our approach within the Crowdsourcing-LLM intersection
In this paper, we integrate a broad survey of literature on LLM chaining and crowdsourcing workflows into a design space. We then use the case studies to explore open questions in this design space. Our broad investigation and resulting design space inform a comprehensive set of recommendations for navigating design decisions and orienting future work. These recommendations span from higher-level guidance, such as adhering to the tradeoff between outcome quality and resource constraints, to lower-level suggestions, such as pursuing architecture-based methods for quality validation. These recommendations expand upon related work by several other groups investigating the intersection of crowdsourcing and LLMs. One such paper explores the impact on requester and crowdworker stakeholders of incorporating LLMs into crowdsourcing workflows (ALLEN et al., 2023). A second considers hybrid Crowd-LLM data annotation (Li, 2024). Another implements three strategies from crowdsourcing using LLMs for data processing tasks (Parameswaran et al., 2023). Closest to our work, the final paper has students implement six crowdsourcing workflows. The paper discusses the effectiveness of these workflows and the impact of the difference between LLM and crowdworker abilities (Wu et al., 2023b). We present a wider scope of recommendations informed by the design space and our case studies.
Title of ”Objectives, what the designer wants the workflow to accomplish”. Three rows of text each with two objectives, examples, and a description. Row 1. A tradeoff between outcome quality (ex. high factuality, creativity) and resource constraints (ex. effort, latency, expense). Description: ”Designers must work within resource constraints to improve outcome quality”. Row 2. A spectrum between objectivity (ex. math, question-answering) and subjectivity (ex. writing, prototyping). Description: Different tasks may require objectivity, subjectivity, or some combination thereof. Row 3. A spectrum between task specific (ex. only applied to one problem (e.g., translation workflow)) and task general (ex. can be applied to many problems(e.g., map-reduce)). Description: Workflows vary from supporting a specific task to supporting a general set of tasks.
@@ 3. Design space methodology
We survey papers that involve the design and implementation of crowdsourcing workflows or LLM chains. Although crowdsourcing workflows and LLM chains have many overlapping concepts, the terms used to denote these concepts differ. In this paper, we refer to both LLMs and crowdworkers as “workers.” We use the term “workflow” to refer to both crowdsourcing workflows and LLM chains. We refer to intermediate steps of workflows as “subtasks.” We use the term “user” to describe the person receiving workflow outputs (e.g., the “requester” in crowdsourcing terms). The term “designer” in this paper refers to a person who defines a workflow to later be used by a “user.” Throughout the discussion of the design space and strategies, we indicate which references are from the crowdsourcing workflow literature [C] and the LLM chaining literature [L].
@@ 3.1. Paper collection
Workflow literature. We collect papers through a search of the keywords: crowd(sourcing) workflow, crowd(sourcing) pipeline, crowdsourcing complex work and microtask. We search the ACM CHI Conference on Human Factors in Computing Systems, ACM Conference on Computer-supported Cooperative Work and Social Computing, the ACM Symposium on User Interface Software and Technology, and the AAAI Conference on Human Computation and Crowdsourcing, as well as Semantic Scholar (Fricke, 2018) and Google Scholar.
Chaining literature. We collect papers through a search of the keywords: LLM chain(ing), prompt chain(ing), LLM pipeline, and AI chain(s/ing). We search the International Conference on Learning Representations, the Conference on Neural Information Processing Systems, the International Conference on Machine Learning, the Association for Computational Linguistics proceedings, the Association for Computing Machinery proceedings, as well as Semantic Scholar (Fricke, 2018) and Google Scholar. We include preprints in our literature search.
Paper inclusion criteria. We include papers that implement, or support the implementation of, a workflow with at least two steps; i.e., workflows in which the outputs from one worker inform the input to another worker. We include papers that focus on building an interface or domain specific language for other designers to build workflows. We restrict our search to academic papers, rather than software like LangChain (LangChain, 2022). We exclude works that call multiple workers in parallel but never in a chain of steps (e.g., majority voting of labels). This rule limits the crowdsourcing techniques considered to those that reflect the serial and dependent structure of an LLM chain. We also exclude techniques such as Chain of Thought (Wei et al., 2022) which issue multiple tasks to a worker in one step instead of incorporating multiple worker calls (Cheng et al., 2015). We include instances in which workers synchronously work on the same interface, viewing and editing each other’s outputs. Finally, the task at hand must have a user working towards some end goal, so we do not consider open-ended collaborative projects like Wikipedia. We make one exception for a paper that explicitly draws from crowdsourcing strategies to complete data aggregation tasks with LLMs (Parameswaran et al., 2023). Although they do not implement chains, the connection to crowdsourcing resonates with our design space goals.
@@ 3.2. Coding strategy
We conduct a thematic analysis (Braun and Clarke, 2006), following the methodologies of similar design spaces in HCI (Lai et al., 2023; Bae et al., 2022a; Shi et al., 2021). First, one author extracted 1) the outcomes (objectives) and 2) the elements (tactics) of the workflows from the paper set. We then employed iterative open coding on the extracted data, in which two authors separately coded each data extraction, sorted resulting codes into sub-categories, and then discussed these categorizations. In three further rounds of iteration, the same authors refined these categorizations. All authors reviewed the resulting design space.
@@ 4. Design Space
Our design space has two parts: the objectives span the space of a designer’s goals, and the tactics span the space of workflow implementations. These objectives are 1) promoting outcome qualities while respecting resource constraints, 2) specifying an objective or subjective definition of quality, and 3) defining the generality or specificity of the set of tasks that the designer wants the workflow to support. For the workflow tactics, actors complete various subtasks that are linked together into the workflow’s architecture.
Although we have no guarantee that the design space is complete, it is based on an exhaustive survey of existing work. Defining this space enables understanding, classification, and comparison of existing workflows (Figure 4 and Section 4.3), generation of new workflows (Section 6), and the discovery of design patterns (Section 5) that inform chaining practices and guide directions for future work (Section 8).
Title: Tactics, the building blocks of workflows. One row each for actors, subtasks, and architectures, with the category and (a short description). The actor’s row contains: Crowdworkers (people on an online platform), LLMs (generative large language models and equivalents), user (the person deploying the workflow). The subtasks row contains: generate (create content), evaluate (judge content), improve (improve content), focus (select content), partition (split tasks), and merge (combine content). The architecture row contains: sequential (call LLMs in sequence), branching (diverge on paths), redundant (multiple actors repeat subtask), dynamic (Initialize parts of the workflow at runtime), and communicative (facilitate interaction among actors) Each subtask category is connected with a generic example instruction. Generate: ”Write a caption”. Evaluate: ”Is this answer grammatical?”. Improve ”Fix the errors” Focus: ”Select regions for future work: Partition: ”Divide the task into subtasks”. Merge: ”Compose many outlines into one”.
@@ 4.1. Objectives in the design space
The objectives of a workflow design define the designer’s goals within several dimensions. Designers will design for some point along a outcome quality vs resource constraints tradeoff, in a range from objectivity to subjectivity, and in a range spanning between task specificity and generality. (Figure 2).
@@ 4.1.1. Outcome quality vs resource constraints
When working towards outcome qualities, designers must be mindful of resource constraints. Workflows can support quality outputs such as by improving accuracy and/or spurring creativity [C(Lasecki et al., 2013) L(Kim et al., 2023b; Creswell and Shanahan, 2022; Reppert et al., 2023; Wu et al., 2021)], transparency [L(Creswell et al., 2022; Creswell and Shanahan, 2022; Wu et al., 2023b; Reppert et al., 2023; Wu et al., 2023a; Kim et al., 2023b)], debuggability [L(Wu et al., 2022a, 2021)], and modularity [C(LaToza et al., 2014; Lee et al., 2017; Retelny et al., 2014) L(Huang et al., 2023b; Zharikova et al., 2023; Yao et al., 2024; Wu et al., 2023a)].
@@
Workflows can also address, although imperfectly, quality concerns. Some quality concerns are induced by the workers, such as hallucinations [L(Nair et al., 2023a; Mirowski et al., 2023; Creswell and Shanahan, 2022; Nair et al., 2023b; Wu et al., 2023a; Parameswaran et al., 2023)] and generative loops [L(Mirowski et al., 2023)] in the case of LLMs, and fatigue [C(Dai et al., 2013; Alshaibani et al., 2020)] in the case of crowdworkers. Others are induced by the workflow itself, such as cascading errors [C(Kittur et al., 2011; Goto et al., 2016) L(Wu et al., 2022a, 2021; Xie et al., 2023)] and a lack of coherency [C(Bernstein et al., 2010; Kim et al., 2017; Hahn et al., 2016) L(Yang et al., 2022; Mirowski et al., 2023; Wu et al., 2022a)]. More concerns are induced by the broader context, such as the need to scale [C(Kulkarni et al., 2014, 2012) L(Arora et al., 2023)], the workflow’s placement within a larger system [C(Pavel et al., 2014; Cheng and Bernstein, 2015; Christoforaki and Ipeirotis, 2014)], unethical use of the workflow [L(Park et al., 2022)], and the impact on human labor [L(Wu et al., 2023a; Mirowski et al., 2023)], sense of ownership [L(Mirowski et al., 2023)] and sense of agency [L(Mirowski et al., 2023; Kim et al., 2023b)]. However, mechanisms for improving output quality can be more expensive [C(Huffaker et al., 2020) L(Yao et al., 2024)], take longer [C(Chang et al., 2017; Gouravajhala et al., 2018)], and require more effort to create [C(Kittur et al., 2011; LaToza et al., 2014)].
@@
Expense. The expense of a workflow refers to paying crowdworkers or paying for an API call to an LLM. Workflows are often more expensive than simpler approaches (e.g., Chain of Thought prompting) [C(Marcus et al., 2011a, b; Hahn et al., 2016; Lin et al., 2014a; Dai et al., 2010; Huffaker et al., 2020) L(Yao et al., 2024; Wang et al., 2022; Nair et al., 2023a; Du et al., 2023; Gero et al., 2023; Zharikova et al., 2023)]. However, the cost is often less than hiring professionals [C(Liem et al., 2011; Christoforaki and Ipeirotis, 2014; Zaidan and Callison-Burch, 2011; Ambati et al., 2012) L(Fleming et al., 2023)], pretraining models [L(Huang et al., 2023b; Levine et al., 2022; Kim et al., 2023b; Cobbe et al., 2021)], or completing the task yourself [L(Zhang et al., 2022; Wang et al., 2023)].
@@
Latency. Another constraint is the time for the workflow to complete [C(Kulkarni et al., 2014; Lasecki et al., 2015; Marcus et al., 2011b, a)]. Many workflows have lower latency than a single person or a baseline performing the same task, as work can be parallelized across multiple workers [C(Pietrowicz et al., 2013; Kulkarni et al., 2014; Lasecki et al., 2015; Chilton et al., 2014; André et al., 2013; Teevan et al., 2016; Lin et al., 2014a; Gouravajhala et al., 2018; Lee et al., 2017; Guo et al., 2016) L(Kim et al., 2023b)].
@@
Effort. Workflows can help users complete tasks with less effort [C(Kulkarni et al., 2014) L(Wu et al., 2023a; Mirowski et al., 2023; Long et al., 2023)]. However, workflows can require high effort to create [C(Hahn et al., 2016; Kulkarni et al., 2012; Kittur et al., 2012)] and oversee [C(LaToza et al., 2014) L(Wu et al., 2022a)].
4.1.2. Objectivity vs subjectivity